OpenStudy (anonymous):

What is the slope of the line determined by the points (5,-3) and (-9,-6)? A) 3/14 B) -3/14 C) 9/4 D) 14/3

OpenStudy (anonymous):

Is it y=m+b form?

OpenStudy (whpalmer4):

No. you have two points, call them \((x_1,y_1)\) and \((x_2,y_2)\) The slope is given by the formula \[m = \frac{y_2-y_1}{x_2-x_1}\] It doesn't make a bit of difference which point you decide is \((x_1,y_1)\) — you just have to make sure you are consistent, and don't change them around for the bottom half of the fraction!

OpenStudy (anonymous):

oh, ok

OpenStudy (whpalmer4):

I would suggest a convention of using the first point listed as \((x_1,y_1)\) and the second point listed as \((x_2,y_2)\)

OpenStudy (anonymous):

-6-3/-9-5

OpenStudy (anonymous):

9/4

OpenStudy (whpalmer4):

Nope. the first point is (5,-3), NOT (5,3)

OpenStudy (whpalmer4):

They put that answer in there to catch people who aren't paying careful attention to their work.

OpenStudy (anonymous):

oh

OpenStudy (whpalmer4):

The actual equation is \[m = \frac{-6-(-3)}{-9-5}\] Do you see your two mistakes? One of them is in writing the equation, and the other in evaluating it.

OpenStudy (anonymous):

let me do it again

OpenStudy (whpalmer4):

I can't stop you :-)

OpenStudy (anonymous):

I still got 9/4

OpenStudy (anonymous):

-6-(-3)/(-9)-5=9/4

OpenStudy (anonymous):

oh, I add the 9 and 5

OpenStudy (anonymous):

The answer is A

OpenStudy (whpalmer4):

no. what is -6 -(-3)?

OpenStudy (anonymous):

9

OpenStudy (whpalmer4):

let's go through this and brush up your arithmetic with negative numbers. -6 - (-3) = -6 + 3. subtracting a positive number moves you to the left on the number line. subtracting a negative number moves you to the right on the number line.

OpenStudy (anonymous):

I still don't understand why it's not 9

OpenStudy (whpalmer4):

how could it possibly be 9? You aren't starting out with 6, you're starting out with -6. if you owe me $6, and I decide to reduce your debt by $3, do you suddenly have $9 in your pocket?

OpenStudy (anonymous):

oh, ok, I understand now

OpenStudy (whpalmer4):

what is -9 -5?

OpenStudy (anonymous):

14

OpenStudy (whpalmer4):

no. again, if you owe me $9, and I reduce your debt by $5 (I subtract -$5 from your debt), do you suddenly have $14? I don't think so.

OpenStudy (anonymous):

4

OpenStudy (anonymous):

I thought if you have 2 negative numbers, they become positive

OpenStudy (anonymous):

negative signs*

OpenStudy (whpalmer4):

Draw a number line. 

OpenStudy (whpalmer4):

Let's do -3 - 1 Start at -3. Subtracting a positive number (we are subtracting +1) moves you to the left on the number line. -3 - (+1) takes you from -3 1 unit left to -4. -3 -1 = -4

OpenStudy (whpalmer4):

What is -3 -3?

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

So, it's like adding, but with negative signs

OpenStudy (whpalmer4):

-5?!? If it's just like adding, is 3+3 =5 ?

OpenStudy (anonymous):

-6 lol sorry

OpenStudy (whpalmer4):

how old are you, btw?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

I turn 16 in may

OpenStudy (anonymous):

So, the answer is B?

OpenStudy (whpalmer4):

\[m = \frac{-6-(-3)}{-9-5}\] -6-(-3) = -9-5 =

OpenStudy (anonymous):

-3

OpenStudy (whpalmer4):

I asked two questions there...

OpenStudy (anonymous):

-14?

OpenStudy (whpalmer4):

okay, what is -3/-14? is that one of your answer choices, or the equivalent of one of them?

OpenStudy (anonymous):

B?

OpenStudy (whpalmer4):

do you think the - signs are just some sort of decoration that can be added or removed when the fancy strikes? is b) -3/-14?

OpenStudy (anonymous):

So, it's D?

OpenStudy (whpalmer4):

as you typed it above, it appears to be -3/14 -3/14 -3/-14 same, or different?

OpenStudy (anonymous):

diffrent

OpenStudy (whpalmer4):

d) 14/3 14/3 -3/-14 same, or different?

OpenStudy (anonymous):

same

OpenStudy (whpalmer4):

really. 14 divided into 3 equal pieces is the same as -3 divided into -14 equal pieces?

OpenStudy (anonymous):

ugh, this is soooo complicated

OpenStudy (anonymous):

I would say A, but you said that was wrong

OpenStudy (whpalmer4):

no, I didn't say that. what is -3 * -1?

OpenStudy (anonymous):

Yes you did, scroll up

OpenStudy (whpalmer4):

no, I was responding to something else. some of your imaginative arithmetic.

OpenStudy (whpalmer4):

what is -3 * -1?

OpenStudy (anonymous):

-4

OpenStudy (whpalmer4):

do you realize that this symbol: * means multiply?

OpenStudy (anonymous):

right, I new that

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

So, the answer is A?

OpenStudy (whpalmer4):

"knew", not "new" if you knew that meant multiplication, why did you add the two numbers and give me -4 as the result? what does -3*1 =, if -3*-1 = -3?

OpenStudy (anonymous):

3

OpenStudy (whpalmer4):

hmm. what does 2*3 = ?

OpenStudy (whpalmer4):

isn't 2*3 the same as adding 3+3? isn't 3*3 the same as adding 3+3+3? isn't 1*3 the same as just having 3? wouldn't by that same reasoning -3*1 = 1* -3 be the same as just having -3? a positive number * a positive number gives you a positive number a negative number * a positive number gives you a negative number a negative number * a negative number gives you a positive number -3*-1 = 3 -14*-1 = 14 we can multiply the numerator and denominator of a fraction by the same number (as long as it isn't 0) and not change the value at all. \[\frac{-3}{-14} * \frac{-1}{-1} = \frac{3}{14}\] The answer is \[\frac{3}{14}\]

OpenStudy (anonymous):

2x3=6

OpenStudy (anonymous):

I've been saying that for a while now, I said the answer was A

OpenStudy (whpalmer4):

You weren't able to get through the steps to correctly determine that the answer was A. If you had had to write down the answer instead of choosing from a small list, you wouldn't have been able to get the correct answer. That's why I stayed around instead of just saying "okay, good job" and moving on to something else.

OpenStudy (anonymous):

Ok, so maybe I guessed a little...